"Customers arrive at a suburban ticket outlet at the rate of 14 per hour on Monday mornings (exponential interarrival times). Selling the tickets and providing general information takes an average of 3 minutes per customer and varies exponentially. There is 1 ticket agent on duty on Mondays. How many minutes does the average customer spend in the system?"

a). 7

b). 8

c). 9

d). 10

e). 11
Answer #1

The correct answer is A) 7

Given:

Λ mean rate of arrival = 14

µ: mean service rate = 60/3 = 20

Wq : Mean wait in the queue: Λ/ µ(µ- Λ)

Wq => 14/20(20-14) = > 14/120 = > 0.116 => 0.166*60 = 7 Minutes